# Decomposition-Based Optimal Bounds for Privacy Amplification via Shuffling

### v1

@@ 5. New Algorithm on Privacy Amplification
Suppose $W\sim\omega$ is a $\mathbb{Y}$-valued random variable sampled from the blanket. For any $\epsilon>0$ and $x,x^{\prime}\in\mathbb{X}$, the privacy amplification random variable is defined as
$L^{x,x^{\prime}}_{\epsilon}=\frac{R(x)(W)-e^{\epsilon}R(x^{\prime})(W)}{\omega(W)}.$
Using PARV, Balle et al. derived the precise expression for $D_{e^{\epsilon}}(P_{0}^{B}\|P_{1}^{B})$:
Lemma 2 (Lemma 5.3 in (Balle et al., 2019)).
$D_{e^{\epsilon}}$ $(\mathcal{M}_{S}(X^{0})\|\mathcal{M}_{S}(X^{1}))$
(4) $\leq D_{e^{\epsilon}}(P_{0}^{B}\|P_{1}^{B})=\frac{1}{\gamma n}\sum_{m=1}^{n}\binom{n}{m}\gamma^{m}(1-\gamma)^{n-m}\mathbb{E}\left[\sum_{i=1}^{m}L_{i}\right]_{+}.$
The bound above can also be expressed probabilistically as follows (Balle et al., 2019). Let $M\sim\text{Bin}(n,\gamma)$ be the random variable counting the number of users who sample from the blanket of $\mathcal{R}$. Formula (4) can be re-written as:
$D_{e^{\epsilon}}(P_{0}^{B}\|P_{1}^{B})=\frac{1}{\gamma n}\mathbb{E}_{M\sim\text{Bin}(n,\gamma)}\left[\sum_{i=1}^{M}L_{i}\right]_{+},$
where we use the convention $\sum_{i=1}^{m}L_{i}=0$ when $m=0$.
Unfortunately, Balle et al. stopped at this point and did not pursue further simplification. In the following, we demonstrate how to improve upon the PARV to enable precise computation.
Definition 0 (Generalized Privacy Amplification Random Variable (GPARV)).
Define the generalized privacy amplification random variable as:
$\text{G}_{\epsilon}^{x,x^{\prime}}=\begin{cases}\frac{1}{\gamma}L_{\epsilon}^{x,x^{\prime}},&\text{w.p.}\ \gamma,\\ 0,&\text{w.p.}\ 1-\gamma.\end{cases}$
where $L_{\epsilon}^{x,x^{\prime}}$ is defined in Definition 1.
We can now restate Lemma 2 in a simplified form:
@@
$D_{e^{\epsilon}}(\mathcal{M}_{S}(X^{0})\|\mathcal{M}_{S}(X^{1}))\leq D_{e^{\epsilon}}(P_{0}^{B}\|P_{1}^{B})=\frac{1}{n}\mathbb{E}\left[\sum_{i=1}^{n}G_{i}\right]_{+}.$
@@
  1. (1)
    $\mathbb{E}[G]=1-e^{\epsilon}$,
  2. (2)
    $1-e^{\epsilon+\epsilon_{0}}\leq G\leq e^{\epsilon_{0}}-e^{\epsilon}$.
@@
$\mathbb{E}[G]=\gamma\mathbb{E}\left[\frac{1}{\gamma}L\right]=\mathbb{E}[L]=\mathbb{E}_{W\sim\omega}\left[\frac{R(x)(W)-e^{\epsilon}R(x^{\prime})(W)}{\omega(W)}\right]=1-e^{\epsilon}.$
The second property is due to the $\epsilon_{0}$-DP property of $R$: $\forall x\in\mathbb{X},y\in\mathbb{Y}:1\leq\frac{R(x)(y)}{\gamma\omega(y)}\leq e^{\epsilon_{0}}$, so
$1-e^{\epsilon+\epsilon_{0}}\leq\frac{R(x)(W)-e^{\epsilon}R(x^{\prime})(W)}{\gamma\omega(W)}\leq e^{\epsilon_{0}}-e^{\epsilon}.\qed$
@@
In the original paper of the privacy blanket framework, a similar property of PARV was provided, but in a loose form (Balle et al., 2019). Specifically, they established that $\gamma(e^{-\epsilon_{0}}-e^{\epsilon+\epsilon_{0}})\leq L\leq\gamma(e^{\epsilon_{0}}-e^{\epsilon-\epsilon_{0}})$. However, the $\epsilon_{0}$-DP property of $R$ actually guarantees a tighter bound: $\gamma(1-e^{\epsilon+\epsilon_{0}})\leq L\leq\gamma(e^{\epsilon_{0}}-e^{\epsilon})$.
5.2. New Algorithm for Computing Privacy Amplification Upper Bounds
We present a new algorithm for computing the optimal privacy amplification bound under the general clone paradigm for any specific local randomizer, as described in Algorithm 4.
@@ 5.2. New Algorithm for Computing Privacy Amplification Upper Bounds
First, the distribution of the generalized privacy amplification random variable (GPARV) $G_{\epsilon}^{x,x^{\prime}}$ is computed for a given local randomizer. Since most local randomizers exhibit input symmetry, the distribution of $G$ typically does not depend on the specific values of $x$ and $x^{\prime}$. We thus denote it simply by $G_{\epsilon}$. The distributions of $G_{\epsilon}$ for commonly used local randomizers are summarized in Table 1.
Given the distribution of $G_{\epsilon}$, our algorithm discretizes it to obtain $\bar{G}$, by rounding each value in $G$ up to the nearest larger multiple of a discretization interval length $l$:
@@
Because $G$ is supported on the interval $[1-e^{\epsilon+\epsilon_{0}},e^{\epsilon_{0}}-e^{\epsilon}]$ (Property 2), the discretization step requires $O\left(\frac{e^{\epsilon_{0}}}{l}\right)$ operations. For most frequency oracles used on categorical data, $G$ takes values on at most five points (see Table 1), resulting in $O(1)$ discretization time.
@@
The FFT computation runs in $O\left(\frac{n}{l}\log\left(\frac{n}{l}\right)\right)=\widetilde{O}(\frac{n}{l})$ time. Choosing $l=O\left(\frac{1}{n}\right)$ ensures an $O(1)$ additive error in total $\widetilde{O}(n^{2})$ time.
More precise analysis can be done:
@@
where $a=-\mathbb{E}[G]=e^{\epsilon}-1$ and $b=(e^{\epsilon_{0}}-1)(e^{\epsilon}+1)$. This shows that the error decays exponentially in $n$ as long as $l\leq\frac{2a^{2}}{b^{2}}$. Empirical evaluations confirm that $l=O(1)$ is sufficiently accurate in practice (see Section 7), and the overall FFT runtime becomes $O(n\log n)$.
Comparison with Existing Numerical Methods.
The numerical computation of privacy amplification in the shuffle model has been studied since the introduction of the clone paradigm by Feldman et al. (Feldman et al., 2022). Prior numerical algorithms can only handle three-point decompositions, such as computing $D_{e^{\epsilon}}(P_{0}^{C}\|P_{1}^{C})$ (Feldman et al., 2022; Wang et al., 2024; Koskela et al., 2021). However, these techniques do not generalize to decompositions involving more than three points, which are essential for obtaining tight bounds from optimal decompositions.
In contrast, our FFT-based algorithm supports the optimal decomposition of any local randomizer, enabling tighter and more accurate privacy amplification bounds. Furthermore, our method is not only more general but also simpler and significantly faster than existing numerical algorithms (see Section 7).
Table 1. The probability distribution of GPARV for common LDP protocols
$1-e^{\epsilon+\epsilon_{0}}$ $e^{\epsilon_{0}}-e^{\epsilon+\epsilon_{0}}$ $1-e^{\epsilon}$ 0 $e^{\epsilon_{0}}-e^{\epsilon}$
$k$-RR (Warner, 1965) $\frac{1}{e^{\epsilon_{0}}+k-1}$ 0 $\frac{k-2}{e^{\epsilon_{0}}+k-1}$ $\frac{e^{\epsilon_{0}}-1}{e^{\epsilon_{0}}+k-1}$ $\frac{1}{e^{\epsilon_{0}}+k-1}$
BLH (Wang et al., 2017) $\frac{1}{2(e^{\epsilon_{0}}+1)}$ $\frac{1}{2(e^{\epsilon_{0}}+1)}$ $\frac{1}{2(e^{\epsilon_{0}}+1)}$ $\frac{e^{\epsilon_{0}}-1}{e^{\epsilon_{0}}+1}$ $\frac{1}{2(e^{\epsilon_{0}}+1)}$
RAPPOR (Erlingsson et al., 2014) $\frac{1}{(e^{\epsilon_{0}/2}+1)^{2}}$ $\frac{1}{e^{\epsilon_{0}/2}(e^{\epsilon_{0}/2}+1)^{2}}$ $\frac{e^{\epsilon_{0}/2}}{(e^{\epsilon_{0}/2}+1)^{2}}$ $1-e^{-\epsilon_{0}/2}$ $\frac{1}{(e^{\epsilon_{0}/2}+1)^{2}}$
OUE (Wang et al., 2017) $\frac{1}{2(e^{\epsilon_{0}}+1)}$ $\frac{1}{2e^{\epsilon_{0}}(e^{\epsilon_{0}}+1)}$ $\frac{e^{\epsilon_{0}}}{2(e^{\epsilon_{0}}+1)}$ $\frac{1}{2}(1-e^{-\epsilon_{0}})$ $\frac{1}{2(e^{\epsilon_{0}}+1)}$
HR (Acharya et al., 2019) $\frac{1}{2(e^{\epsilon_{0}}+1)}$ $\frac{1}{2(e^{\epsilon_{0}}+1)}$ $\frac{1}{2(e^{\epsilon_{0}}+1)}$ $\frac{e^{\epsilon_{0}}-1}{e^{\epsilon_{0}}+1}$ $\frac{1}{2(e^{\epsilon_{0}}+1)}$
@@
the distritbuion density $G$ of $G_{\epsilon}$, number of users $n$, discretisation interval length $l$
@@
5.3. New Amplification Lower Bounds
Our generalized privacy amplification random variable (GPARV) also facilitates the computation of lower bounds for privacy amplification via shuffling. These bounds help to demonstrate the tightness of the upper bounds.
An upper bound refers to the existence of a value $\delta_{u}$ such that, for a given local randomizer with specified $\epsilon_{0},\epsilon,n$, and for any two neighboring input datasets $X^{0}$ and $X^{1}$, we have:
$D_{e^{\epsilon}}\big{(}\mathcal{M}_{S}(X^{0}),\mathcal{M}_{S}(X^{1})\big{)}\leq\delta_{u}.$
For a given local randomizer with specified $\epsilon_{0},\epsilon,n$, we can construct two neighboring datasets $X^{0}$ and $X^{1}$, and compute:
$\delta_{l}=D_{e^{\epsilon}}\big{(}\mathcal{M}_{S}(X^{0}),\mathcal{M}_{S}(X^{1})\big{)},$
@@ 5.3. New Amplification Lower Bounds
In previous studies, the common approach to selecting $X^{0}$ and $X^{1}$ is to set $x_{2}=x_{3}=\dots=x_{n}$ such that $x_{1}^{0},x_{1}^{1},x_{2}$ are mutually distinct (Feldman et al., 2023a; Wang et al., 2024). The following theorem provides an efficient method for computing $D_{e^{\epsilon}}\big{(}\mathcal{M}_{S}(X^{0}),\mathcal{M}_{S}(X^{1})\big{)}$ under this setting.
@@
Let $\mathcal{R}:\mathbb{X}\to\mathbb{Y}$ be a local randomizer, and let $\mathcal{M}_{S}=\mathcal{S}\circ\mathcal{R}^{n}$ be the shuffling of $\mathcal{R}$. Fix $\varepsilon\geq 0$ and inputs $X^{0}\simeq X^{1}$ with $x_{1}^{0}\neq x_{1}^{1}$ and $x_{2}=x_{3}=\dots=x_{n}$. Define a random variable $G=\frac{\mathcal{R}(x_{1}^{0})(y)-e^{\epsilon}\mathcal{R}(x_{1}^{1})(y)}{\mathcal{R}(x_{2})(y)}$, where $y\sim\mathcal{R}(x_{2})$. Suppose $G_{1},G_{2},\dots$ are i.i.d. copies of $G$. Then, we have the following:
$D_{e^{\epsilon}}(\mathcal{M}_{S}(X^{0})\|\mathcal{M}_{S}(X^{1}))=\frac{1}{n}\mathbb{E}\left[\sum_{i=1}^{n}G_{i}\right]_{+}.$
We provide the proof in Appendix B. The result indicates that, for $X^{0}$ and $X^{1}$ satisfying $x_{2}=x_{3}=\dots=x_{n}$, we can efficiently compute $D_{e^{\epsilon}}\big{(}\mathcal{M}_{S}(X^{0}),\mathcal{M}_{S}(X^{1})\big{)}$ using FFT.
6. Joint Composition of LDP
We assume that any two input vectors $\boldsymbol{x}=(x_{1},x_{2},\dots,x_{m})$ and $\boldsymbol{x}^{\prime}=(x_{1}^{\prime},x_{2}^{\prime},\dots,x_{m}^{\prime})$ are neighboring. In practice, it is common to use the joint composition of $m$ mechanisms, each satisfying $\left(\frac{\epsilon}{m}\right)$-LDP, to achieve overall $\epsilon$-LDP.
Theorem 1 (Composition Theorem).
If $\mathcal{R}_{i}$ satisfies $\epsilon_{i}$-LDP, then $\mathcal{R}^{\times}_{[1:m]}$ satisfies $\sum_{i=1}^{m}\epsilon_{i}$-LDP.
When applied to LDP protocols resulting from joint composition, the clone paradigm provides particularly loose decompositions. This is due to the presence of many intermediate states in the joint probability distribution after the composition. For instance, most local randomizers have probability ratios between any two inputs that belong to the set $\{e^{-\epsilon},1,e^{\epsilon}\}$. However, after a $m$-fold joint composition, the ratio of the joint probability distribution may belong to $\{e^{\frac{i}{m}\epsilon}|i\in[-m,m]\cap\mathbb{Z}\}$. Under these conditions, the common part in the clone paradigm deviates significantly from the actual privacy blanket.
To compute the privacy amplification induced by the privacy blanket, we only need to compute the GPARV of $\mathcal{R}^{\times}_{[1:m]}$, which is equivalent to finding the optimal decomposition of $\mathcal{R}^{\times}_{[1:m]}$. Fortunately, the optimal decomposition of a joint composition mechanism is simply the Cartesian product of the optimal decompositions of each individual LDP component.
Example 6.0 (Optimal Decomposition of Two-Joint $k$-RR).
$\mathcal{R}(a_{1}^{0},b_{1}^{0})=\left(e^{\epsilon_{0}/2}p\mathbbm{1}_{a_{1}^{0}}+p\mathbbm{1}_{a_{1}^{1}}+qU\right)\times\left(e^{\epsilon_{0}/2}p\mathbbm{1}_{b_{1}^{0}}+p\mathbbm{1}_{b_{1}^{1}}+qU\right),$
$\mathcal{R}(a_{1}^{1},b_{1}^{1})=\left(p\mathbbm{1}_{a_{1}^{0}}+e^{\epsilon_{0}/2}p\mathbbm{1}_{a_{1}^{1}}+qU\right)\times\left(p\mathbbm{1}_{b_{1}^{0}}+e^{\epsilon_{0}/2}p\mathbbm{1}_{b_{1}^{1}}+qU\right),$
$\forall i\in[2,n],\ \mathcal{R}(a_{i},b_{i})=\left(p\mathbbm{1}_{a_{1}^{0}}+p\mathbbm{1}_{a_{1}^{1}}+qU+(1-2p-q)\mathbbm{1}_{a_{i}}\right)$
×$\quad\times\left(p\mathbbm{1}_{b_{1}^{0}}+p\mathbbm{1}_{b_{1}^{1}}+qU+(1-2p-q)\mathbbm{1}_{b_{i}}\right)$
where $p=\frac{1}{e^{\epsilon_{0}/2}+k-1}$, $q=\frac{k-2}{e^{\epsilon_{0}/2}+k-1}$, and $U$ is the uniform distribution over $[k]-\{a_{1}^{0},a_{1}^{1}\}$. The Cartesian product of two probability distributions $P,Q$ is defined as $(P\times Q)(a,b)=P(a)\cdot Q(b)$.
7. Numerical Experiments
In this section, we conduct experimental evaluations of our FFT-based numerical algorithm proposed in this paper.

### v5

@@ 5.1 Generalized privacy amplification random variable
Suppose $W\sim\omega$ is a $\mathbb{Y}$-valued random variable sampled from the blanket. For any $\varepsilon>0$ and $x,x^{\prime}\in\mathbb{X}$, Define two random variable:
$L^{x,x^{\prime}}_{\varepsilon}$ $=\frac{\mathcal{R}(x)(W)-e^{\varepsilon}\mathcal{R}(x^{\prime})(W)}{\omega(W)}.$
$\text{G}_{\varepsilon}^{x,x^{\prime}}$ =$=\begin{cases}\frac{1}{\gamma}L_{\varepsilon}^{x,x^{\prime}},&\text{w.p.}\ \gamma,\\ 0,&\text{w.p.}\ 1-\gamma.\end{cases}$
The quantity $L_{\varepsilon}^{x,x^{\prime}}$ was introduced by Balle et al. [BBGN19] and referred to as the Privacy Amplification Random Variable (PARV). We refer to $G_{\varepsilon}^{x,x^{\prime}}$ as the Generalized Privacy Amplification Random Variable (GPARV). The following theorem follows directly from Theorem 12 and the definition of the primary optimal decomposition in Section 4.2.
@@
$D_{e^{\varepsilon}}(\mathcal{M}_{S}(X^{0})\parallel\mathcal{M}_{S}(X^{1}))\leq D_{e^{\varepsilon}}(P_{0}^{B}\parallel P_{1}^{B})=\frac{1}{n}\mathbb{E}\left[\sum_{i=1}^{n}G_{i}\right]_{+}.$
We now briefly revisit how the original privacy blanket paper utilizes the PARV. Using PARV, Balle et al. derived the precise expression for $D_{e^{\varepsilon}}(P_{0}^{B}\parallel P_{1}^{B})$:
Lemma 3 (Lemma 5.3 in [BBGN19]).
$D_{e^{\varepsilon}}(\mathcal{M}_{S}(X^{0})\parallel\mathcal{M}_{S}(X^{1}))\leq D_{e^{\varepsilon}}(P_{0}^{B}\parallel P_{1}^{B})$ $=\frac{1}{\gamma n}\sum_{m=1}^{n}\binom{n}{m}\gamma^{m}(1-\gamma)^{n-m}\mathbb{E}\left[\sum_{i=1}^{m}L_{i}\right]_{+}$
$=\frac{1}{\gamma n}\mathbb{E}_{M\sim\text{Bin}(n,\gamma)}\left[\sum_{i=1}^{M}L_{i}\right]_{+},$ (11)
where $M\sim\text{Bin}(n,\gamma)$. We use the convention $\sum_{i=1}^{m}L_{i}=0$ when $m=0$.
Unfortunately, Balle et al. stopped at this point and did not pursue further simplification. In fact, the proof of Theorem 13 can be obtained by algebraically manipulating the expression in Lemma 3, using the definitions of PARV and GPARV. However, a more fundamental derivation arises directly from Theorem 12 and the definition of the general clone reduction with the primary optimal decomposition. By introducing the general clone framework, we provide a simpler and more principled proof of the specific bound established by the privacy blanket.
From the perspective of GPARV, the construction of the simplified optimal decomposition in Section 4.4 becomes quite natural. The reason why components of the primary optimal decomposition that share the same ratio $a_{j}:b_{j}:c_{j}$ can be merged is that the corresponding GPARV takes the same value on each of these components, namely $\frac{a_{j}-e^{\varepsilon}b_{j}}{c_{j}}$. It can be verified that the GPARV defined by the decomposition remains unchanged after such merging.
@@
  1. 1.
    $\mathbb{E}[G]=1-e^{\varepsilon}$,
  2. 2.
    $1-e^{\varepsilon+\varepsilon_{0}}\leq G\leq e^{\varepsilon_{0}}-e^{\varepsilon}$.
@@
$\mathbb{E}[G]=\gamma\mathbb{E}\left[\frac{1}{\gamma}L\right]=\mathbb{E}[L]=\mathbb{E}_{W\sim\omega}\left[\frac{\mathcal{R}(x)(W)-e^{\varepsilon}\mathcal{R}(x^{\prime})(W)}{\omega(W)}\right]=1-e^{\varepsilon}.$
The second property is due to the $\varepsilon_{0}$-DP property of $\mathcal{R}$: $\forall x\in\mathbb{X},y\in\mathbb{Y}:1\leq\frac{\mathcal{R}(x)(y)}{\gamma\omega(y)}\leq e^{\varepsilon_{0}}$, so
$1-e^{\varepsilon+\varepsilon_{0}}\leq\frac{\mathcal{R}(x)(W)-e^{\varepsilon}\mathcal{R}(x^{\prime})(W)}{\gamma\omega(W)}\leq e^{\varepsilon_{0}}-e^{\varepsilon}.$
@@
In the original paper of the privacy blanket framework, a similar property of PARV was provided, but in a loose form [BBGN19]. Specifically, they established that $\gamma(e^{-\varepsilon_{0}}-e^{\varepsilon+\varepsilon_{0}})\leq L\leq\gamma(e^{\varepsilon_{0}}-e^{\varepsilon-\varepsilon_{0}})$. However, the $\varepsilon_{0}$-DP property of $\mathcal{R}$ actually guarantees a tighter bound: $\gamma(1-e^{\varepsilon+\varepsilon_{0}})\leq L\leq\gamma(e^{\varepsilon_{0}}-e^{\varepsilon})$.
5.2 Algorithm for computing privacy amplification upper bounds
We present a new algorithm for computing the optimal privacy amplification upper bound under the general clone paradigm for any specific local randomizer, as described in Algorithm 7.
@@ 5.2 Algorithm for computing privacy amplification upper bounds
First, the distribution of the generalized privacy amplification random variable $G_{\varepsilon}^{x,x^{\prime}}$ is computed for a given local randomizer. Since many local randomizers exhibit input symmetry, the distribution of $G$ typically does not depend on the specific values of $x$ and $x^{\prime}$. We thus denote it simply by $G_{\varepsilon}$. For local randomizers lacking input symmetry, additional considerations are required; we discuss this case in detail in Section 8.
The distributions of $G_{\varepsilon}$ for commonly used local randomizers are summarized in Table 1. Since the optimal decompositions and corresponding GPARVs of BLH, RAPPOR, and OUE depend on the domain size $D$ (see Table 2 in Appendix A), we present their GPARVs in the asymptotic regime (i.e., for $D\gg 1$) for simplicity. The distribution of $G_{\varepsilon}$ for Laplace mechanism on $\{0,1\}$ is presented in Appendix B. (The Laplace mechanism over the interval $[0,1]$ does not satisfy input symmetry.)
Given the distribution of $G_{\varepsilon}$, our algorithm discretizes it to obtain $\bar{G}$, by rounding each value in $G$ up to the nearest larger multiple of a discretization interval length $l$:
@@
Because $G$ is supported on the interval $[1-e^{\varepsilon_{0}+\varepsilon},e^{\varepsilon_{0}}-e^{\varepsilon}]$ (Property 2), the discretization step requires $O\left(\frac{e^{\varepsilon_{0}+\varepsilon}}{l}\right)$ operations. For most frequency oracles used on categorical data, $G$ takes values on at most five points (see Table 1), resulting in $O(1)$ discretization time.
@@
The FFT computation runs in $O\left(\frac{n}{l}\log\left(\frac{n}{l}\right)\right)=\widetilde{O}(\frac{n}{l})$ time. Choosing $l=O\left(\frac{1}{n}\right)$ ensures an $O(1)$ additive error in total $\widetilde{O}(n^{2})$ time. More precise analysis can be done:
@@
where $a=-\mathbb{E}[G]=e^{\varepsilon}-1$ and $b=(e^{\varepsilon_{0}}-1)(e^{\varepsilon}+1)$. This shows that the error decays exponentially in $n$ as long as $l\leq\frac{2a^{2}}{b^{2}}$.
The above analysis concerns the error introduced when using the FFT-based algorithm to compute $D_{e^{\varepsilon}}\big{(}P_{0}^{B}\parallel P_{1}^{B}\big{)}$. From a theoretical perspective, since $\delta$ is typically very small, achieving high-precision computation (i.e., with error $o(\delta)$) would require a very small discretization step size $l$, which incurs a high computational cost.
However, our experimental results show that when discretization is used to infer the smallest value of $\varepsilon$ such that $D_{e^{\varepsilon}}\big{(}P_{0}^{B}\parallel P_{1}^{B}\big{)}\leq\delta$, the resulting error in $\varepsilon$ is negligible in practice. This is because the distribution of the GPARV is highly sensitive to changes in $\varepsilon$. In particular, once $\varepsilon$ exceeds the true threshold $\varepsilon^{*}$ (defined by $D_{e^{\varepsilon^{*}}}\big{(}P_{0}^{B}\parallel P_{1}^{B}\big{)}=\delta$), the function $\delta(\varepsilon)$ decreases rapidly. This rapid decay ensures that even a relatively coarse discretization yields an accurate estimate of $\varepsilon$. Empirical evaluations confirm that $l=O(1)$ is sufficiently accurate in practice (see Section 6), and the overall FFT runtime becomes $O(n\log n)$.
@@
the distritbuion density $G$ of $G_{\varepsilon}$, number of users $n$, discretisation interval length $l$
@@
Table 1: The probability distribution of GPARV for common local randomizers
(The GPARVs for BLH, RAPPOR and OUE are asymptotic as $|\mathbb{X}|\to\infty$.) $1-e^{\varepsilon_{0}+\varepsilon}$ $e^{\varepsilon_{0}}-e^{\varepsilon_{0}+\varepsilon}$ $1-e^{\varepsilon}$ 0 $e^{\varepsilon_{0}}-e^{\varepsilon}$ $k$-RR [War65] $\frac{1}{e^{\varepsilon_{0}}+k-1}$ 0 $\frac{k-2}{e^{\varepsilon_{0}}+k-1}$ $\frac{e^{\varepsilon_{0}}-1}{e^{\varepsilon_{0}}+k-1}$ $\frac{1}{e^{\varepsilon_{0}}+k-1}$ BLH [WBLJ17] $\frac{1}{2(e^{\varepsilon_{0}}+1)}$ $\frac{1}{2(e^{\varepsilon_{0}}+1)}$ $\frac{1}{2(e^{\varepsilon_{0}}+1)}$ $\frac{e^{\varepsilon_{0}}-1}{e^{\varepsilon_{0}}+1}$ $\frac{1}{2(e^{\varepsilon_{0}}+1)}$ RAPPOR [EPK14] $\frac{1}{(e^{\varepsilon_{0}/2}+1)^{2}}$ $\frac{1}{e^{\varepsilon_{0}/2}(e^{\varepsilon_{0}/2}+1)^{2}}$ $\frac{e^{\varepsilon_{0}/2}}{(e^{\varepsilon_{0}/2}+1)^{2}}$ $1-e^{-\varepsilon_{0}/2}$ $\frac{1}{(e^{\varepsilon_{0}/2}+1)^{2}}$ OUE [WBLJ17] $\frac{1}{2(e^{\varepsilon_{0}}+1)}$ $\frac{1}{2e^{\varepsilon_{0}}(e^{\varepsilon_{0}}+1)}$ $\frac{e^{\varepsilon_{0}}}{2(e^{\varepsilon_{0}}+1)}$ $\frac{1}{2}(1-e^{-\varepsilon_{0}})$ $\frac{1}{2(e^{\varepsilon_{0}}+1)}$ HR [ASZ19] $\frac{1}{2(e^{\varepsilon_{0}}+1)}$ $\frac{1}{2(e^{\varepsilon_{0}}+1)}$ $\frac{1}{2(e^{\varepsilon_{0}}+1)}$ $\frac{e^{\varepsilon_{0}}-1}{e^{\varepsilon_{0}}+1}$ $\frac{1}{2(e^{\varepsilon_{0}}+1)}$
Comparison with Existing Numerical Methods.
The numerical computation of privacy amplification in the shuffle model has been studied since the introduction of the clone paradigm by Feldman et al. [FMT22]. Prior numerical algorithms can only handle three-point decompositions, such as computing $D_{e^{\varepsilon}}(P_{0}^{C}\parallel P_{1}^{C})$ [FMT22, WPL+24]. However, these techniques do not generalize to decompositions involving more than three points, which are essential for obtaining tight bounds from optimal decompositions.
In contrast, our FFT-based algorithm supports the optimal decomposition of any local randomizer, enabling tighter and more accurate privacy amplification bounds. Furthermore, our method is not only more general but also simpler and faster than existing numerical algorithms (see Section 6).
5.3 Algorithm for computing privacy amplification lower bounds
An upper bound refers to the existence of a value $\delta_{u}$ such that, for a given local randomizer with specified $\varepsilon_{0},\varepsilon,n$, and for any two neighboring input datasets $X^{0}$ and $X^{1}$, we have:
$D_{e^{\varepsilon}}\big{(}\mathcal{M}_{S}(X^{0})\parallel\mathcal{M}_{S}(X^{1})\big{)}\leq\delta_{u}.$
For a given local randomizer with specified $\varepsilon_{0},\varepsilon,n$, we can construct two neighboring datasets $X^{0}$ and $X^{1}$, and compute:
$\delta_{l}\leq D_{e^{\varepsilon}}\big{(}\mathcal{M}_{S}(X^{0})\parallel\mathcal{M}_{S}(X^{1})\big{)},$
@@ 5.3 Algorithm for computing privacy amplification lower bounds
In previous studies, a common strategy for selecting neighboring datasets $X^{0}$ and $X^{1}$ is to set $x_{2}=x_{3}=\dots=x_{n}$, such that $x_{1}^{0}$, $x_{1}^{1}$, and $x_{2}$ are mutually distinct [FMT23a, WPL+24, GKMP20]. In the special case where $|\mathbb{X}|=2$, the datasets $X^{0}$ and $X^{1}$ can be chosen as
$x_{1}^{0}\neq x_{1}^{1},\quad\text{and}\quad x_{2}=x_{3}=\dots=x_{n}=x_{1}^{1}.$
For the Laplace mechanism on $\mathbb{X}=\{0,1\}$, we choose
$x_{1}^{0}=0,\quad x_{1}^{1}=1,\quad\text{and}\quad x_{2}=x_{3}=\dots=x_{n}=1.$
The following theorem provides an efficient method for computing the hockey-stick divergence $D_{e^{\varepsilon}}\big{(}\mathcal{M}_{S}(X^{0})\parallel\mathcal{M}_{S}(X^{1})\big{)}$ under this setting. It appeared in [GKMP20].
@@
Let $\mathcal{R}:\mathbb{X}\to\mathbb{Y}$ be a local randomizer, and let $\mathcal{M}_{S}=\mathcal{S}\circ\mathcal{R}^{n}$ be the shuffling of $\mathcal{R}$. Fix $\varepsilon\geq 0$ and inputs $X^{0}\simeq X^{1}$ with $x_{1}^{0}\neq x_{1}^{1}$ and $x_{2}=x_{3}=\dots=x_{n}$. Define a random variable $G=\frac{\mathcal{R}(x_{1}^{0})(y)-e^{\varepsilon}\mathcal{R}(x_{1}^{1})(y)}{\mathcal{R}(x_{2})(y)}$, where $y\sim\mathcal{R}(x_{2})$. Suppose $G_{1},G_{2},\dots$ are i.i.d. copies of $G$. Then, we have the following:
$D_{e^{\varepsilon}}(\mathcal{M}_{S}(X^{0})\parallel\mathcal{M}_{S}(X^{1}))=\frac{1}{n}\mathbb{E}\left[\sum_{i=1}^{n}G_{i}\right]_{+}.$
Proof.
It is a direct result from Theorem 12 by setting $A_{0}=\mathcal{R}(x_{1}^{0}),A_{1}=\mathcal{R}(x_{1}^{1})$ and $A_{2}=\mathcal{R}(x_{2})$. ∎
Therefore, we can compute the lower bound using an algorithm similar to Algorithm 7, with the only difference being that rounding up is replaced by rounding down. This modification ensures the correctness of the resulting bound. For convenience, we refer to the random variable $G$ in Theorem 14 as the lower-bound GPARV.
Theorem 14 also provides a method for computing the lower bound in the cases of joint composition and parallel composition. The procedure is analogous to that used for computing the corresponding upper bound: one first computes the decomposition of each sub-local randomizer at inputs $x_{1}^{0},x_{1}^{1},x_{2}$. Since $x_{2}=x_{3}=\cdots=x_{n}$, the resulting decomposition does not involve any left-over distribution. Each component of the decomposition is defined over the set $\{y\mid\mathcal{R}_{i}(x_{1}^{0})(y):\mathcal{R}_{i}(x_{1}^{1})(y):\mathcal{R}_{i}(x_{2})(y)=a:b:c\}$.
By applying either the Cartesian product (for joint composition) or a weighted combination (for parallel composition) of these components, one obtains the decomposition corresponding to the composed local randomizer. This decomposition can then be used to compute the lower-bound GPARV.
It is important to note that, for joint composition, only the decompositions of the sub-local randomizers satisfy the compositional structure—this does not necessarily extend to their lower-bound GPARV. In contrast, for parallel composition, the lower-bound GPARV can be obtained as a weighted mixture of the individual lower-bound GPARVs for the sub-local randomizers.
Comparison with existing lower bound methods:
While [GKMP20] provided Theorem 14, they explicitly note that their computational method is applicable only when the induced random variable $G$ has support of size three. For example, consider $k$-RR ($k\geq 4$) with datasets $X=(1,3,\dots,3)$ and $X^{\prime}=(2,3,\dots,3)$. The random variable
$\frac{\mathcal{R}(x_{1}^{0})(y)-e^{\varepsilon}\mathcal{R}(x_{1}^{1})(y)}{\mathcal{R}(x_{2})(y)},\quad y\sim\mathcal{R}(x_{2})$
[…]
6 Numerical Experiments
In this section, we evaluate the performance of our FFT-based numerical algorithm for computing optimal privacy amplification bounds under the general clone framework.
Comparison with existing methods.